OpenStudy (anonymous):

What is the square root of 252 in simplified radical form?

OpenStudy (alexandervonhumboldt2):

sqrt252 ghere the biggest perfect square number is 36. so sqrt(36*7)=sqrt(6*6*7)=\[6\times \sqrt{7}\]

OpenStudy (anonymous):

could u check my other answers please

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

is the \[\sqrt{28}\] approximated to the nearest tenth? 5.3

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Is the \[\sqrt{144}\] approximated to the nearest tenth? 10.7

OpenStudy (alexandervonhumboldt2):

no this is a perfect square number sqrt144=12